OpenStudy (anonymous):

Find the exact value of the expression cos75 degrees sin15 degrees+sin75 degrees cos 15degrees.

OpenStudy (cwrw238):

use the identity sin ( a + b) = sin a cos b + cos a sin b

OpenStudy (anonymous):

what identity is

OpenStudy (anonymous):

that exactly>?

OpenStudy (cwrw238):

if we take a to be 75, and b to be 15 then sin (a + b) = sin (75 + 15) = sin 90 and what is the value of sin 90?

OpenStudy (anonymous):

soorrhy but i dnt understand this is there an

OpenStudy (anonymous):

another way?

OpenStudy (cwrw238):

i dont think so look at the formula again and read it from left to right

OpenStudy (cwrw238):

sin a cos b + cos a sin b = sin ( a + b) so sin 75 cos 15 + cos 75 sin 15 = sin ( 75 + 15) (replacing a by 75 and b by 15)

OpenStudy (cwrw238):

and sin 90 = 1 so the exact value is 1